OpenStudy (anonymous):

the average of two numbers a and b is given by A = a + b / 2 and the average of two numbers is 75 and one of the numbers is 91 then what is the other number solve and show work

OpenStudy (agreene):

\[A= \frac{a+b}2\] with A given as 75 and a given as 91 \[75=\frac{91+b}2\] \[150= 91+b\] \[59 = b\]